a) Find f'(x) using logarithmic differentiation, where f(x) = e−3x√2x−5 (6−5x)4
.
(d) Evaluate the integralZ(x3 + 1)1/3x5dx.
1
Expert's answer
2020-07-01T17:20:23-0400

a. f(x) = exp(-3"x(2x)" 0.5 "- 5(6-5x)4" )

Taking natural logarithm on both sides:

"ln f(x) = -3x(2x)" 0.5 - "20(6-5x)"

or, "ln f(x) = -3x(2x)" 0.5 - "120 + 100x"

"=-3(2)" 0.5"x" (1.5) "- 120 + 100x"

Differentiating w.r.t "x:" "-(1\/f(x)) * d(f(x))\/dx" "= -3*(2)" 0.5 * "x" (1.5+1)"\/(1.5+1)" "+100"

or, "(1\/f(x))*d(f(x))\/dx=( -3*(2)" 0.5 * ("x)" (2.5) )*("2\/5)" +"100"

or, "d(f(x))\/dx =" "[(-6*2" 0.5 * "x" 2.5)"\/5 + 100]" * exponent("-3x * (" "2x)" 0.5 "-20[6-5x]" (Answer)


b) "Z= \int" "(x" 3"+1)\/(3x^5)" "dx"

"= \int[x^3/(3x^5) + 1/(3x^5)]dx"

"=\int[1/(3x^2) + 1/(3x^5)]dx"

"=(1/3) \int[ 1/x^2+ 1/x^5]dx"

"=(1\/3) [x" (-2+1)/(-2+1) + "x" (-5+1)/(-5+1)] + C

"=(1\/3)[-(1\/x) - 1\/(4x^4)] + c"

"=-(1\/3)[(1\/x) + 1\/(4x^4)] +c" (Answer)